OpenStudy (anonymous):

y varies jointly as x and z. Write the appropriate joint-variation equation, and find y for the given values of x and z. y = -108 when x = -4 and z = 3; x = 6 and z = -2 a. y = 9xz; -108 c. y = -6xz; -104 b. y = -7xz; 392 d. y = ; 60

OpenStudy (anonymous):

Again, joint variation means y is directly proportional to both x and z. I.e. y = kx and y = kz, to combine = y = kxz, for some constant k. 1) Finding the constant k by solving for it simultaneously. y = -108 when x = -4 and z = 3; x = 6 and z = -2 -108 = k (-4)(3) k = ? With this given k, substitute inside the below equation to find y. y = k (6)(-2)

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

Yes